• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Seonho Choi

    Log entry time 22:53:53 on December13,2007

    Entry number 220544

    keyword=Configuration Change

    Both spectrometers have moved to 60 degrees. (Actually 60.003 for left and 60.001 degrees for right.) New momentum setting at 713.329 MeV/c. Both Q2/Q3 were recycled (degaussed). L-HRS is ready, but R-HRS dipole is still ramping to the target value. Start to take data with L-HRS.